# Twinkle Twinkle Little Star, Roman Sees Where You Are: Predicting Exoplanet Transit Yields in the Rosette Nebula with the Nancy Grace Roman Space Telescope

### v1

@@ Abstract
Young stars host only a small fraction of the known exoplanet population because their photometric variability, magnetic activity, and frequent placement in dense, poorly-resolved regions hamper exoplanet detections. Yet, measuring planets at these ages is crucial since these phases are when dynamical processes that drive planetary migration are most active. We assess the expected yield of a hypothetical Nancy Grace Roman Space Telescope transit survey of the Rosette Nebula, a $\sim 10\,\mathrm{Myr}$ star-forming region with a dense and diverse stellar population. Using the Roman Exposure Time Calculator to quantify sensitivity to Rosette members, we establish detection thresholds for companions and evaluate yields via Monte Carlo injection-recovery simulations. We predict the detection of $33\pm 9$ young transiting exoplanets orbiting stellar hosts in a month-long survey, and $29\pm 8$ in a two-week survey. The extended baseline primarily improves sensitivity to longer-period planets orbiting FGK stars, while most M dwarf detections are well-sampled within two weeks. Irrespective of the temporal baseline, transit detections are dominated by of 1-2 $R_{\oplus}$ super-Earths and sub-Neptunes with $P\lesssim 8$ days. Such a sample would substantially expand the census of only three detected planets younger than $20\,\mathrm{Myr}$, probing an age regime in which planetary radii remain inflated, the stability of close-in orbits is uncertain, and planetary migration may still be ongoing. This survey offers a path to constrain early planetary evolution and establish prime follow-up targets for the James Webb Space Telescope, Vera Rubin Observatory, and the Habitable Worlds Observatory.
@@
facilities: Gaia, James Webb Space Telescope, Mikulski Archive for Space Telescopes, NASA Exoplanet Archivesoftware: Astroquery (A. Ginsburg et al., 2019), Astropy (Astropy Collaboration et al., 2013, 2022; V. Almendros-Abad et al., 2023), batman (L. Kreidberg, 2015), dustmaps (G. M. Green, 2018), HTCondor (D. Thain et al., 2005), Lightcurve (Lightkurve Collaboration et al., 2018) MIST (A. Dotter, 2016; J. Choi et al., 2016; B. Paxton et al., 2011, 2013, 2015), NumPy (C. R. Harris et al., 2020), PAdova and TRieste Stellar Evolution Code (A. Bressan et al., 2012; Y. Chen et al., 2014), pandas (W. McKinney, 2010), pysynphot (STScI Development Team, 2013), scikit-learn (F. Pedregosa et al., 2011), SciPy (P. Virtanen et al., 2020)
@@ 1 Introduction
Understanding how planetary systems form and evolve during their first few million years remains one of the central challenges in exoplanetary science. During this early epoch, planets are still contracting and cooling (J. J. Fortney et al., 2007; E. F. Linder et al., 2019), while their orbits and atmospheres are being actively sculpted by disk-driven migration, dynamical interactions, and stellar evolution (G. M. Kennedy & S. J. Kenyon, 2008; S. Ida & D. N. C. Lin, 2008; C. Cossou et al., 2014). Detecting and characterizing young systems in this stage therefore provides a direct window into the physical mechanisms that determine planetary radii, compositions, and orbital architectures. These observations can help discriminate between in-situ and migration-dominated formation pathways (C. Mordasini et al., 2012) and constrain the timescales for planets to reach thermal and dynamical equilibria (J. E. Owen, 2019). Furthermore, this assessment is also critical to understand planetary habitability, as the locations of habitable zones also evolve rapidly during the first few hundred million years (R. K. Kopparapu et al., 2013), influencing the composition of forming planets (V. S. Meadows et al., 2018). Observing planets over this formative stage allows us to probe the progenitor population of mature, potentially habitable worlds.
@@
However, young transiting planets remain rare. There are only a small number of confirmed planets younger that $\lessapprox 100$ Myr (e.g., A. W. Mann et al., 2017; T. J. David et al., 2019, 2016; T. Holczer et al., 2016; M. G. Barber et al., 2024). Observations of planets at distinct formative stages are necessary to trace the evolutionary pathways of these systems. Nevertheless, strong optical extinction and crowded environments in young star-forming regions (SFRs) make such observations extremely challenging with current ground-based and space observatories.
@@
The Nancy Grace Roman Space Telescope (hereafter Roman), NASA’s next flagship astrophysics mission, is designed to deliver wide-field imaging and spectroscopy and provides an unprecedented opportunity to overcome these limitations. Scheduled to launch in late 2026, it will carry the Wide Field Instrument (WFI), a 300 megapixel near-infrared camera comprised of 18 detectors, and a $0.28\,\mathrm{deg}^{2}$ field of view (FOV), approximately 100 times larger than the Hubble Space Telescope’s FOV. Observations at these near-infrared wavelengths are less affected by dust extinction relative to optical bands ($A_{\mathrm{J}}/A_{\mathrm{V}}=0.282$, $A_{\mathrm{H}}/A_{\mathrm{V}}=0.175$, and $A_{\mathrm{K}}/A_{\mathrm{V}}=0.112$), enabling a more complete census of reddened and embedded sources. Furthermore, unlike previous optical wide-field transit surveys such as Kepler, TESS, HATNet, and WASP, which become incomplete for low-mass hosts ($\lesssim 0.2\;M_{\odot}$) due to limited sensitivity at faint optical magnitudes (W. J. Borucki et al., 2010; G. R. Ricker et al., 2015; G. A. Bakos, 2018; A. M. S. Smith et al., 2006), observing in the near-infrared allows detections around cooler and redder low-mass stars.
@@
Several recent studies have demonstrated Roman’s potential for exoplanet discovery across distinct Galactic environments: the Transiting Exosatellites, Moons, and Planets in Orion (TEMPO) survey (TEMPO I: M. A. Limbach et al. 2023; TEMPO II: M. Soares-Furtado et al. 2024) established its capability to detect and characterize transiting planets, satellites, and exomoons in the Orion Nebula Cluster (ONC) with a month-long survey; J. C. Wilson et al. (2019) estimate that Roman’s Galactic Bulge Time-Domain Survey could detect up to $\sim 10^{5}$ transiting exoplanets; S. K. Terry et al. (2025) constrain exoplanet detections with microlensing events; and E. Kerins et al. (2023) outline the proposed Roman Survey of the Earth Transit Zone (RoSETZ) to search for habitable-zone planets around nearby stars. Collectively, these efforts illustrate Roman’s transformative capability to map the architectures of planetary systems across a wide range of stellar ages and environments.
@@
Building on this foundation, we outline a potential investigation of the Rosette Nebula, a dense SFR with a rich population of slightly older and more massive stars (Section 3), in comparison to the ONC. This complementary environment offers a unique laboratory for investigating the formation pathways of planetary systems during the late disk-dissipation phase (E. Gaidos et al., 2025).
@@
This paper is organized as follows: Section 2 details our survey design elements such as our target field of the Rosette Nebula, FOV orientation, and filter selection. In Section 3, we present membership analysis and determination of this region’s (sub)stellar population. We present the expected signal-to-noise for objects in the Rosette Nebula in Section 4. In Section 5, we introduce and discuss our modeling methodology to determine exoplanet transit yields. Finally, in Section 6, we provide a brief summary of this work.
2 Observational Setup
@@
2.1 Target Field Selection
There exists a large population of young and dense SFRs in the Milky Way, including the ONC (1–3 Myr; R. D. Jeffries, 2007), Upper Scorpius, Lagoon Nebula (1–2 Myr; L. Venuti et al., 2021), 30 Doradus (1–2 Myr; J. Sun et al., 2022), the Rosette Nebula (0.5–10 Myr; K. Mužić et al., 2019), and Westerhout 40 (<7<7 Myr; R. Y. Shuping et al., 2012). Among these, the Rosette Nebula stands out as an especially compelling target for a Roman exoplanet transit survey because (1) the region hosts thousands of young stellar objects spanning a wide range of spectral types (O5–L2) (e.g., L. Cambrésy et al., 2013; K. Mužić et al., 2019), (2) its distance (1400 pc) places members in an ideal regime across a range of (sub)stellar masses that is bright enough for high S/N yet faint enough to avoid saturation, and (3) the region comfortably fits within Roman’s FOV as shown in Figure 1, with an estimated peak surface density of $\Sigma=300\mathrm{\;stars\;pc^{-2}}$ (M. A. Kuhn et al., 2014, 2015).
Figure 1: Roman WFI footprint overlaid on an image of the Rosette Nebula taken from the Second Digitized Sky Survey (DSS-2; B. M. Lasker et al., 1990). Red squares outline the 18 individual WFI detectors. The physical scale bar assumes a distance of 1.4 kpc to the Rosette Nebula.
At the system’s distance of approximately 1400 pc (K. Mužić et al., 2019), FGK-type stars remain below the Roman Wide Field Instrument’s saturation limit, while high photometric precision (Table 2) is preserved for low-mass stars and brown dwarfs (BDs). Moreover, Gaia (Gaia Collaboration et al., 2016) astrometric measurements reveal that the stars in the Rosette Nebula occupy a single, well-defined locus (see middle panel in Figure 2), indicating largely aligned kinematic motions. This property simplifies membership analysis of the region in comparison to kinematically complex regions with multiple velocity components such as Westerhout 40 (T. Shimoikura et al., 2015).
@@ 2.1 Target Field Selection
2.2 Field of View & Orientation
We defined the survey footprint by combining the Gaia DR3 catalog, including sources down to a faintness limit of $G\approx 21$ mag (A. Vallenari et al., 2023), with the Roman WFI footprint available through the Aladin Lite v3 interactive sky atlas (M. Baumann et al., 2022). [^1] https://aladin.cds.unistra.fr/AladinLite/ For this analysis, we considered sources falling within the boundaries of the 18 WFI detectors, omitting dithers. To explore the impact of orientation, we sampled 12 different position angles in 30 increments. The field center was set to coincide with the densest stellar region at the heart of the Rosette Nebula, located at $(\alpha,\delta)=(06^{\mathrm{h}}\,31^{\mathrm{m}}\,45.81^{\mathrm{s}},+04^{\circ}\,51^{\prime}\,58.0^{\prime\prime})$, with an orientation angle of $270^{\circ}$ selected to maximize the number of monitored stars. In this orientation, the Roman footprint encompasses a total of 11,352 stars from the Gaia DR3 catalog. We use this as our precursory sample to determine the stellar population that Roman will monitor in the Rosette Nebula (as discussed in Section 3).
2.3 Filter Selection
@@
2.4 Exposure Time and Cadence Selection
Similar to the exposure time and cadence selection protocol of the TEMPO survey, adopting 6-read integrations with individual exposure times of 18 seconds. This configuration ensures the minimum number of integrations recommended for both high-precision photometry and effective cosmic-ray mitigation, while ensuring that we preserve the dynamic high signal-to-noise across bright and faint stellar populations. We also note that our resulting cadence is sufficient to temporally resolve ingress and egress, including close-in planets whose shortest simulated transit (Section 5) lasts $\sim 2$ hr.
3 Assessing the (Sub)Stellar Population of the Rosette Nebula
@@ 3.1 Membership Determination
We began with the 11,352 Gaia DR3 sources that overlap the Roman/WFI footprint shown as a combination of gray and salmon-hued points in the left panel of Figure 2. To reduce foreground and background contamination, we applied a distance prior restricting sources to 1–1.9 kpc. This range was guided by previous distance estimates for the Rosette Nebula (e.g., K. Mužić et al., 2019, 2022), but expanded to be slightly more generous in order to allow for the possibility of additional members at larger separations.
@@
To identify a core member of the Rosette Nebula that could serve as a kinematic reference to find comoving member stars, we applied Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN; R. J. G. B. Campello et al., 2013) to five parameters (RA, Dec, proper-motion in RA ($\mu_{\alpha}$), proper-motion in Dec ($\mu_{\delta}$), and parallax ($\pi$)) from this sample. The core member Gaia DR3 3131334834450613376 ($\overline{\mathrm{RV}}=38.01\,\mathrm{km\,s^{-1}}$, A. Almeida et al. 2023) was identified. The proper motion of this core member ($\mu_{\alpha}=-1.67\;\mathrm{mas\,yr^{-1}}$, $\mu_{\delta}=0.24\;\mathrm{mas\,yr^{-1}}$) was in good agreement with the proper-motion distribution identified by K. Mužić et al. (2022) in a prior analysis of the Rosette Nebula.
To further refine our sample of likely Rosette Nebula members, we conducted a volume-limited (200 pc) kinematic search using the FriendFinder [^2] https://github.com/adamkraus/Comove algorithm (B. M. Tofflemire et al., 2021) to identify stars whose tangential velocities were within $\pm 10\;\mathrm{km\,s^{-1}}$ of the values predicted for their positions if they were comoving with our chosen core member. We adopted a comparatively generous $\pm 10\;\mathrm{km\,s^{-1}}$ window to avoid excluding bonafide members given the non-virial, expanding kinematics of the Rosette Nebula, where the tangential velocity is expected to increase with radius from the core (B. Lim et al., 2021; K. Mužić et al., 2022). We opted not to include radial velocities (RVs) as an additional kinematic constraint, as most stars in this highly-extincted region lack reliable RV measurements.
This procedure yielded 546 candidate Rosette Nebula members. As illustrated in Figure 2, the candidate members are more concentrated toward the center of the Roman/WFI footprint, following the known Rosette Nebula structure, while the field stars are spread more uniformly across the field. Further, in proper-motion space, the Rosette Nebula members form a tight clump with mean values of $\bar{\mu}_{\alpha}=-1.58\pm 0.5\,\mathrm{mas\,yr^{-1}}$ and $\bar{\mu}_{\delta}=0.23\pm 0.5\,\mathrm{mas\,yr^{-1}}$, which is distinct from the broader distribution of field stars and overlap the distributions from prior investigations (K. Mužić et al., 2022). In distance space, the Rosette Nebula members exhibit a strong peak centered at 1.4 kpc (see Figure 2), which we adopt as our distance to the Rosette Nebula in further analyses. Our adopted distance is consistent with literature estimates for the Rosette Nebula: 1.39 kpc (H. Hensberge et al., 2000), 1.48 kpc (K. Mužić et al., 2022), 1.59 kpc (K. Mužić et al., 2019), and 1.6 kpc (R. L. Phelps & J. E. Ybarra, 2005).

### v2

@@ Abstract
Young stars host only a small fraction of the known exoplanet population because their photometric variability, magnetic activity, and frequent placement in dense, poorly-resolved regions hamper exoplanet detections. Yet, measuring planets at these ages is crucial since these phases are when dynamical processes that drive planetary migration are most active. We assess the expected yield of a hypothetical Nancy Grace Roman Space Telescope transit survey of the Rosette Nebula, a ${\sim}10\,\mathrm{Myr}$ star-forming region with a dense and diverse stellar population. Using the Roman Exposure Time Calculator to quantify sensitivity to Rosette members, we establish detection thresholds for companions and evaluate yields via Monte Carlo injection-recovery simulations, accounting for nebular extinction and youth-driven stellar variability. We predict the detection of $33\pm 9$ young transiting exoplanets orbiting stellar hosts in a month-long survey, and $29\pm 8$ in a two-week survey. The extended baseline primarily improves sensitivity to longer-period planets orbiting FGK stars, while most M dwarf detections are well-sampled within two weeks. Irrespective of the temporal baseline, transit detections are dominated by of 1-2 $R_{\oplus}$ super-Earths and sub-Neptunes with $P\lesssim 8$ days. Such a sample would substantially expand the census of only three detected transiting planets younger than $20\,\mathrm{Myr}$ around stars less massive than the Sun, probing an age regime in which planetary radii remain inflated, the stability of close-in orbits is uncertain, and planetary migration may still be ongoing. This survey offers a path to constrain early planetary evolution and establish prime follow-up targets for JWST, Rubin, and the Habitable Worlds Observatory.
@@
facilities: Gaia, James Webb Space Telescope, Mikulski Archive for Space Telescopes, NASA Exoplanet Archivesoftware: Astroquery (Ginsburg et al., 2019), Astropy (Astropy Collaboration et al., 2013, 2022; Almendros-Abad et al., 2023), batman (Kreidberg, 2015), dustmaps (Green, 2018), HTCondor (Thain et al., 2005), Lightcurve (Lightkurve Collaboration et al., 2018) MIST (Dotter, 2016; Choi et al., 2016; Paxton et al., 2011, 2013, 2015), NumPy (Harris et al., 2020), PAdova and TRieste Stellar Evolution Code (Bressan et al., 2012; Chen et al., 2014), pandas (McKinney, 2010), pysynphot (STScI Development Team, 2013), scikit-learn (Pedregosa et al., 2011), SciPy (Virtanen et al., 2020)
@@ I Introduction
Understanding how planetary systems form and evolve during their first few million years remains one of the central challenges in exoplanetary science. During this early epoch, planets are still contracting and cooling (Fortney et al., 2007; Linder et al., 2019), while their orbits and atmospheres are being actively sculpted by disk-driven migration, dynamical interactions, and stellar evolution (Kennedy and Kenyon, 2008; Ida and Lin, 2008; Cossou et al., 2014). Detecting and characterizing young systems in this stage therefore provides a direct window into the physical mechanisms that determine planetary radii, compositions, and orbital architectures. These observations can help discriminate between in-situ and migration-dominated formation pathways (Mordasini et al., 2012) and constrain the timescales for planets to reach thermal and dynamical equilibria (Owen, 2019). Furthermore, this assessment is also critical to understand planetary habitability, as the locations of habitable zones also evolve rapidly during the first few hundred million years (Kopparapu et al., 2013), influencing the composition of forming planets (Meadows et al., 2018). Observing planets over this formative stage allows us to probe the progenitor population of mature, potentially habitable worlds.
@@
However, young transiting planets remain rare. There are only a small number of confirmed planets younger that $\lessapprox 100$ Myr (e.g., Mann et al., 2017; David et al., 2019, 2016; Holczer et al., 2016; Barber et al., 2024). Observations of planets at distinct formative stages are necessary to trace the evolutionary pathways of these systems. Nevertheless, strong optical extinction and crowded environments in young star-forming regions (SFRs) make such observations extremely challenging with current ground-based and space observatories.
@@
The Nancy Grace Roman Space Telescope (hereafter Roman), NASA’s next flagship astrophysics mission, is designed to deliver wide-field imaging and spectroscopy and provides an unprecedented opportunity to overcome these limitations. Scheduled to launch in late 2026, it will carry the Wide Field Instrument (WFI), a 300 megapixel near-infrared camera comprised of 18 detectors, and a $0.28\,\mathrm{deg}^{2}$ field of view (FOV), approximately 100 times larger than the Hubble Space Telescope’s FOV. Observations at these near-infrared wavelengths are less affected by dust extinction relative to optical bands ($A_{\mathrm{J}}/A_{\mathrm{V}}=0.282$, $A_{\mathrm{H}}/A_{\mathrm{V}}=0.175$, and $A_{\mathrm{K}}/A_{\mathrm{V}}=0.112$), enabling a more complete census of reddened and embedded sources. Furthermore, unlike previous optical wide-field transit surveys such as Kepler, TESS, HATNet, and WASP, which become incomplete for low-mass hosts ($\lesssim 0.2\;M_{\odot}$) due to limited sensitivity at faint optical magnitudes (Borucki et al., 2010; Ricker et al., 2015; Bakos, 2018; Smith et al., 2006), observing in the near-infrared allows detections around cooler and redder low-mass stars.
@@
Several recent studies have demonstrated Roman’s potential for exoplanet discovery across distinct Galactic environments: the Transiting Exosatellites, Moons, and Planets in Orion (TEMPO) survey (TEMPO I: Limbach et al. 2023; TEMPO II: Soares-Furtado et al. 2024) established its capability to detect and characterize transiting planets, satellites, and exomoons in the Orion Nebula Cluster (ONC) with a month-long survey; Wilson et al. (2019) estimate that Roman’s Galactic Bulge Time-Domain Survey could detect up to ${\sim}10^{5}$ transiting exoplanets; Terry et al. (2025) constrain exoplanet detections with microlensing events; and Kerins et al. (2023) outline the proposed Roman Survey of the Earth Transit Zone (RoSETZ) to search for habitable-zone planets around nearby stars. Collectively, these efforts illustrate Roman’s transformative capability to map the architectures of planetary systems across a wide range of stellar ages and environments.
@@
Building on this foundation, we outline a potential investigation of the Rosette Nebula, a dense SFR with a rich population of slightly older and more massive stars (Section 3), in comparison to the ONC. This complementary environment offers a unique laboratory for investigating the formation pathways of planetary systems during the late disk-dissipation phase (Gaidos et al., 2025).
@@
This paper is organized as follows: Section II details our survey design elements such as our target field of the Rosette Nebula, FOV orientation, and filter selection. In Section III, we present membership analysis and determination of this region’s (sub)stellar population. We present the expected signal-to-noise for objects in the Rosette Nebula in Section IV. In Section V, we introduce and discuss our modeling methodology to determine exoplanet transit yields. In Section VI, we provide a brief summary of this work. Lastly, in the Appendix we present Kernel Density Estimates (KDEs) of our recovered planet population from these simulations.
II Observational Setup
@@
II.1 Target Field Selection
There exists a large population of young and dense SFRs in the Milky Way, including the ONC (1–3 Myr; Jeffries, 2007), Upper Scorpius, Lagoon Nebula (1–2 Myr; Venuti et al., 2021), 30 Doradus (1–2 Myr; Sun et al., 2022), the Rosette Nebula (0.5–10 Myr; Mužić et al., 2019), and Westerhout 40 (<7<7 Myr; Shuping et al., 2012). Among these, the Rosette Nebula stands out as an especially compelling target for a Roman exoplanet transit survey because (1) the region hosts thousands of young stellar objects spanning a wide range of spectral types (O5–L2) (e.g., Cambrésy et al., 2013; Mužić et al., 2019), (2) its distance (1400 pc) places members in an ideal regime across a range of (sub)stellar masses that is bright enough for high S/N yet faint enough to avoid saturation, and (3) the region comfortably fits within Roman’s FOV as shown in Figure 1, with an estimated peak surface density of $\Sigma=300\mathrm{\;stars\;pc^{-2}}$ (Kuhn et al., 2014, 2015).
Figure 1: Roman WFI footprint overlaid on an image of the Rosette Nebula taken from the Second Digitized Sky Survey (DSS-2; Lasker et al., 1990). Red squares outline the 18 individual WFI detectors. The physical scale bar assumes a distance of 1.4 kpc to the Rosette Nebula.
At the system’s distance of approximately 1400 pc (Mužić et al., 2019), FGK-type stars remain below the Roman Wide Field Instrument’s saturation limit, while high photometric precision (Table 2) is preserved for low-mass stars and brown dwarfs (BDs). Moreover, Gaia (Gaia Collaboration et al., 2016) astrometric measurements reveal that the stars in the Rosette Nebula occupy a single, well-defined locus (see middle panel in Figure 2), indicating largely aligned kinematic motions. This property simplifies membership analysis of the region in comparison to kinematically complex regions with multiple velocity components such as Westerhout 40 (Shimoikura et al., 2015).
@@ II.1 Target Field Selection
II.2 Field of View & Orientation
We defined the survey footprint by combining the Gaia DR3 catalog, including sources down to a faintness limit of $G\approx 21$ mag (Vallenari et al., 2023), with the Roman WFI footprint available through the Aladin Lite v3 interactive sky atlas (Baumann et al., 2022). [^1] https://aladin.cds.unistra.fr/AladinLite/ For this analysis, we considered sources falling within the boundaries of the 18 WFI detectors, omitting dithers. To explore the impact of orientation, we sampled 12 different position angles in 30 increments. The field center was set to coincide with the densest stellar region at the heart of the Rosette Nebula, located at $(\alpha,\delta)=(06^{\mathrm{h}}\,31^{\mathrm{m}}\,45.81^{\mathrm{s}},+04^{\circ}\,51^{\prime}\,58.0^{\prime\prime})$, with an orientation angle of $270^{\circ}$ selected to maximize the number of monitored stars. In this orientation, the Roman footprint encompasses a total of 11,352 stars from the Gaia DR3 catalog. We use this as our precursory sample to determine the stellar population that Roman will monitor in the Rosette Nebula (as discussed in Section III).
II.3 Filter Selection
@@
II.4 Exposure Time and Cadence Selection
Similar to the exposure time and cadence selection protocol of the TEMPO survey, adopting 6-read integrations with individual exposure times of 18 seconds. This configuration ensures the minimum number of integrations recommended for both high-precision photometry and effective cosmic-ray mitigation, while ensuring that we preserve the dynamic high signal-to-noise across bright and faint stellar populations. We also note that our resulting cadence is sufficient to temporally resolve ingress and egress, including close-in planets whose shortest simulated transit (Section V) lasts ${\sim}2$ hr.
III Assessing the (Sub)Stellar Population of the Rosette Nebula
@@ III.1 Membership Determination
We began with the 11,352 Gaia DR3 sources that overlap the Roman/WFI footprint shown as a combination of gray and salmon-hued points in the left panel of Figure 2. To reduce foreground and background contamination, we applied a distance prior restricting sources to 1–1.9 kpc. This range was guided by previous distance estimates for the Rosette Nebula (e.g., Mužić et al., 2019, 2022), but expanded to be slightly more generous in order to allow for the possibility of additional members at larger separations.
@@
To identify a core member of the Rosette Nebula that could serve as a kinematic reference to find comoving member stars, we applied Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN; Campello et al., 2013) to five parameters (RA, Dec, proper-motion in RA ($\mu_{\alpha}$), proper-motion in Dec ($\mu_{\delta}$), and parallax ($\pi$)) from this sample. The core member Gaia DR3 3131334834450613376 ($\overline{\mathrm{RV}}=38.01\,\mathrm{km\,s^{-1}}$, Almeida et al. 2023) was identified. The proper motion of this core member ($\mu_{\alpha}=-1.67\;\mathrm{mas\,yr^{-1}}$, $\mu_{\delta}=0.24\;\mathrm{mas\,yr^{-1}}$) was in good agreement with the proper-motion distribution identified by Mužić et al. (2022) in a prior analysis of the Rosette Nebula.
To further refine our sample of likely Rosette Nebula members, we conducted a volume-limited (200 pc) kinematic search using the FriendFinder [^2] https://github.com/adamkraus/Comove algorithm (Tofflemire et al., 2021) to identify stars whose tangential velocities were within $\pm 10\;\mathrm{km\,s^{-1}}$ of the values predicted for their positions if they were comoving with our chosen core member. We adopted a comparatively generous $\pm 10\;\mathrm{km\,s^{-1}}$ window to avoid excluding bonafide members given the non-virial, expanding kinematics of the Rosette Nebula, where the tangential velocity is expected to increase with radius from the core (Lim et al., 2021; Mužić et al., 2022). We opted not to include radial velocities (RVs) as an additional kinematic constraint, as most stars in this highly-extincted region lack reliable RV measurements.
This procedure yielded 546 candidate Rosette Nebula members. As illustrated in Figure 2, the candidate members are more concentrated toward the center of the Roman/WFI footprint, following the known Rosette Nebula structure, while the field stars are spread more uniformly across the field. Further, in proper-motion space, the Rosette Nebula members form a tight clump with mean values of $\bar{\mu}_{\alpha}=-1.58\pm 0.5\,\mathrm{mas\,yr^{-1}}$ and $\bar{\mu}_{\delta}=0.23\pm 0.5\,\mathrm{mas\,yr^{-1}}$, which is distinct from the broader distribution of field stars and overlap the distributions from prior investigations (Mužić et al., 2022). In distance space, the Rosette Nebula members exhibit a strong peak centered at 1.4 kpc (see Figure 2), which we adopt as our distance to the Rosette Nebula in further analyses. Our adopted distance is consistent with literature estimates for the Rosette Nebula: 1.39 kpc (Hensberge et al., 2000), 1.48 kpc (Mužić et al., 2022), 1.59 kpc (Mužić et al., 2019), and 1.6 kpc (Phelps and Ybarra, 2005).